OpenStudy (anonymous):

dimensional analysis: d = 1/2 vi *t + a*t

OpenStudy (shane_b):

Using SI unites: d = distance (meters or m) vi = velocity (meters/second or m/s) t = time (seconds or s) a = acceleration (meters/second^2 or m/s^2) \[d = \frac{1}{2}V_i *t + a*t \]\[m=\frac{1}{2}\frac{m}{\cancel s}(\cancel s)+\frac{m}{\cancel {s^2}s}(\cancel s)\] The dimensional analysis above shows that the result should be in meters...which is correct.

OpenStudy (anonymous):

what about vf^2 = 2ad? I got m/s = 2kg/s^2m

OpenStudy (shane_b):

I actually don't think I showed the cancellation correctly above when I did all the latex formatting but it does actually work out. Anyway, the formula you posted should be:\[V_f^2-Vi^2=2ad\]\[(\frac{m}{s})^2-(\frac{m}{s})^2=\frac{m}{s^2}m=\frac{m^2}{s^2}=(\frac{m}{s})^2\]

OpenStudy (anonymous):

no, it says vf^2 = 2 ad. but i think it's missing an initial velocity anyways.

OpenStudy (shane_b):

I guess if Vi=0 then it would be correct to simplify it that way

OpenStudy (anonymous):

yea... i think i'm just going to keep it that way lol. not my mistake. but that is true if vi = 0.

OpenStudy (anonymous):

d - vi*t + a*t; i have m = m/1 + m/s. is that correct?

OpenStudy (shane_b):

It looks like you're missing and "=" sign there...

OpenStudy (shane_b):

*an

OpenStudy (anonymous):

You need t^2 for the units to work out correctly: d = v1t +at^2 m = m/s (s) + m/s^2 (s^2)

OpenStudy (shane_b):

That too :)